# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ 7 Discussion
Additional experimental results to complement Section 6 are presented here. Legend names are as in Section 6.
S8.1 Synthetic data
Noisy predictions.
For this experiment, we set $t^{*}=500$ (see Section S7.3) for all methods.
Figure S4 reports the average cumulative miscoverage rate for the results shown in Figure 1. As desired, the cumulative miscoverage rate lies below the threshold $\alpha$ for all $n$.
@@
Figure S5 shows the performance of the Bayes-assisted prediction-powered AsympCS procedures under a Gaussian prior on the noisy predictions experiment in Section 6.1. The results are consistent with those presented in Section 6.1.
@@
Figure S5: Noisy predictions study with Gaussian prior. The left, middle and right panels show average interval volume and cumulative miscoverage rate over $1000$ repetitions for noise levels $\sigma_{Y}=0.1,0.8,3.0$. Results for unassisted PPI++ are shown for reference.
@@
In particular, also with Bayes-assistance, PPI++ easily adapts to increasing noise levels, while standard PPI fails to do so. Moreover, in this case, Bayes-assisted PPI++ outperforms the unassisted version across all noise levels. This is due to the fact that, in this experiment, the predictions from $f$, while noisy, are unbiased for all values of $\sigma_{Y}$. As a result of this, the zero-mean prior used by the Bayes-assisted procedures is well specified, and any additional shrinkage performed by the latter is beneficial.
Biased predictions.
@@
For this experiment, we set $t^{*}=100$ (see Section S7.3) for all methods shown in Figures 2 and S6. The values of $t^{*}$ used for Figure S7 are reported in the figure legend.
@@
Figure S6 reports the average cumulative miscoverage rate for the results shown in Figure 2 at $\upsilon=0$. As discussed in Section 6.1, the cumulative miscoverage rate increases slightly as we decrease $\mathrm{df}$, but remains below the threshold $\alpha$ for all $n$, as desired.
@@
Figure S6: Biased predictions study. The left, middle and right panels show average interval volume and cumulative miscoverage rate over $100$ repetitions for $\mathrm{df}=5,10,\infty$.
@@
Figure S7 repeats the results of Figure S6 for different values of $t^{*}$, which affects the procedure-specific hyperparameters as discussed in Section S7.3. For unassisted PPI, $t^{*}$ represents the time at which the procedure’s interval width is minimised. Therefore, as expected, increasing $t^{*}$ above $100$ leads to larger intervals for $n=100$. However, the qualitative behaviour of the unassisted methods as $\upsilon$ varies is the same across all values of $t^{*}$: their volume remain constant across bias levels, reflecting the lack of prior information. On the other hand, for Bayes-assisted methods, a larger $t^{*}$ implies a smaller prior scale $\tau$. As a result of this, increasing $t^{*}$ above $100$ leads to stronger prior influence at $n=100$. In particular, a large $t^{*}$ results in slightly smaller intervals for $\upsilon\approx 0$, but larger intervals for $|\upsilon|\gg 0$. When comparing the results across different priors, the results are consistent with those in Figure 2: the interval volume under heavier-tailed priors, such as the Laplace and Student-t priors, grow at a lower rate with $|\upsilon|$ compared to the Gaussian prior, thereby offering greater robustness to prior mispecification.
@@
Figure S7: Biased predictions study with different hyperparameters. Each column correspond to one of the prediction-powered methods in Figure 2 for $t^{*}=100,500,1000$. The top and bottom column show average interval volume and cumulative miscoverage rate over $100$ repetitions with $\mathrm{df}=\infty$.
S8.2 Real data
@@
For each of the real data experiments, we set $t^{*}$ (see Section S7.3) equal to the largest $n$ considered in the experiment. In particular, for the flights, forest, and galaxies datasets, we set $t^{*}=10000$, $500$, and $1000$, respectively.
Figure S8 adds the results of the standard PPI procedures to the ones shown in Figure 3.
@@
Figure S8: Real data study. The top and bottom rows show the average interval volume and cumulative miscoverage rate over $1000$ repetitions for the flights, forest, and galaxies datasets.
@@
For these experiments, the improvement of PPI++ over standard PPI is small, and the results remain consistent with those in Figure 3. That is, PPI methods consistently improve over classical inference, with Bayes-assisted methods providing an additional efficiency boost for moderate labelled sample sizes.
Appendix S9 Extended Ville’s inequality and improper prior
As mentioned in Section 3.1, the sequence of intervals $(\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,t}(\overline{Y}_{t},\sigma;\rho))_{t\geq 1}$ defined in Equation 5 constitutes an exact CS for the mean $\mu$ of an i.i.d. sequence of Gaussian random variables with variance $\sigma^{2}$. This follows from combining the method of mixtures for nonnegative martingales, using a local Gaussian prior $\mu\sim\mathcal{N}(\mu;\mu_{0},\rho^{2})$ as mixing density, together with Ville’s inequality (Wang and Ramdas, 2023, Proposition 5.6). That is, the unassisted CS (5) may alternatively be interpreted as a Bayes-assisted CS under a local Gaussian prior, which is non-informative. Crucially, such a construction via Ville’s inequality relies on the fact that the local Gaussian prior is proper, which implies that the associated likelihood ratio, as defined in Equation S26, is a nonnegative martingale. In particular, the same procedure cannot be used to construct a valid CS under the improper prior $\mu\sim(2\pi)^{-1/2}$, which is another popular non-informative prior, but whose associated likelihood ratio does not satisfy the standard martingale integrability condition.
Wang and Ramdas (2023) recently derived an extension of Ville’s inequality that applies to extended nonnegative martingales (Wang and Ramdas, 2023, Definition 3.1), such as the likelihood ratio under the improper prior. In particular, combining the method of mixtures with the extended Ville’s inequality (Wang and Ramdas, 2023, Theorem 4.1) enables to construct a valid CS for the mean $\mu$ of an i.i.d. sequence of Gaussian random variables with variance $\sigma^{2}$ under the improper prior $\mu\sim(2\pi)^{-1/2}$ (Wang and Ramdas, 2023, Corollary 5.9). From this, the proofs of Theorems 3.3 and 3.4 can be easily adapted to the improper prior case, in turn allowing its application as an alternative non-informative prior for unassisted PPI and PPI++ AsympCS (Proposition 5.1).
Below, we repeat the examples presented in Section 6 using the improper prior $\mu\sim(2\pi)^{-1/2}$. When compared to the local Gaussian prior, the prediction-powered AsympCS under the improper prior generally perform slightly worse, at least for the choice of the hyperparameter $\rho$ (see Section S7.3) used here. Indeed, the improper prior may represent a compelling choice in cases where choosing $\rho$ proves problematic, as it does not depend on any hyperparameter.
@@
Figure S9: Noisy predictions study with improper prior. The left, middle and right panels show average interval volume and cumulative miscoverage rate over $1000$ repetitions for noise levels $\sigma_{Y}=0.1,0.8,3.0$. Results for unassisted PPI++ (local Gaussian prior) are shown for reference.
@@
Figure S10: Biased predictions study with improper prior. The left, middle and right panels show average interval volume and cumulative miscoverage rate over $100$ repetitions for $\mathrm{df}=5,10,\infty$. Results for unassisted PPI (local Gaussian prior) are shown for reference.
@@
Figure S11: Real data study with improper prior. The top and bottom rows show the average interval volume and cumulative miscoverage rate over $1000$ repetitions for the flights, forest, and galaxies datasets. Results for classical inference and unassisted PPI++ (local Gaussian prior) are shown for reference.

### v2

@@ 7 Discussion
Additional experimental results to complement Section˜6 are presented here. Legend names are as in Section˜6.
S7.1 Synthetic data
S7.1.1 Noisy predictions
For this experiment, we set $t^{\star}=500$ (see Section˜S6.4) for all methods.
Figure˜S5 reports the average cumulative miscoverage rate for the results shown in Figure˜1. As desired, the cumulative miscoverage rate lies below the threshold $\alpha$ for all $n$.
@@
Figure˜S6 shows the performance of the Bayes-assisted prediction-powered AsympCS procedures under a Gaussian prior on the noisy predictions experiment in Section˜6.1. The results are consistent with those presented in Section˜6.1.
@@
Figure S6: Noisy predictions study with Gaussian prior. The left, middle and right panels show average interval volume and cumulative miscoverage rate over $1000$ repetitions for noise levels $\sigma_{Y}=0.1,0.8,3.0$. Results for non-assisted PPI++ are shown for reference.
@@
In particular, also with Bayes-assistance, PPI++ easily adapts to increasing noise levels, while standard PPI fails to do so. Moreover, in this case, Bayes-assisted PPI++ outperforms the non-assisted version across all noise levels. This is due to the fact that, in this experiment, the predictions from $f$, while noisy, are unbiased for all values of $\sigma_{Y}$. As a result of this, the zero-mean prior used by the Bayes-assisted procedures is well specified, and any additional shrinkage performed by the latter is beneficial.
S7.1.2 Biased predictions
@@
For this experiment, we set $t^{\star}=500$ (see Section˜S6.4) for all methods shown in Figures˜2 and S7. The values of $t^{\star}$ used for Figure˜S8 are reported in the figure legend.
@@
Figure˜S7 reports the average cumulative miscoverage rate for the results shown in Figure˜2 at $\upsilon=0$. As discussed in Section˜6.1, the cumulative miscoverage rate increases slightly as we decrease $\mathrm{df}$, but remains below the threshold $\alpha$ for all $n$, as desired.
@@
Figure S7: Biased predictions study. The left, middle and right panels show average interval volume and cumulative miscoverage rate over $1000$ repetitions for $\mathrm{df}=5,10,\infty$.
@@
Figure˜S8 repeats the simulation of Figure˜S7 for different values of $t^{\star}$, which affects the procedure-specific hyperparameters as discussed in Section˜S6.4. For non-assisted PPI, $t^{\star}$ represents the time at which the procedure’s interval width is minimised. Therefore, as expected, increasing $t^{\star}$ above $100$ leads to larger intervals for $n=100$. However, the qualitative behaviour of the non-assisted methods as $\upsilon$ varies is the same across all values of $t^{\star}$: their volume remain constant across bias levels, reflecting the lack of prior information. On the other hand, for Bayes-assisted methods, a larger $t^{\star}$ implies a smaller prior scale $\tau$. As a result of this, increasing $t^{\star}$ above $100$ leads to stronger prior influence at $n=100$. In particular, a large $t^{\star}$ results in slightly smaller intervals for $\upsilon\approx 0$, but larger intervals for $|\upsilon|\gg 0$. When comparing the results across different priors, the results are consistent with those in Figure˜2: the interval volume under heavier-tailed priors, such as the Laplace and Student-t priors, grow at a lower rate with $|\upsilon|$ compared to the Gaussian prior, thereby offering greater robustness to prior misspecification.
@@
Figure S8: Biased predictions study with different hyperparameters. Each column correspond to one of the prediction-powered methods in Figure˜2 for $t^{\star}=100,500,1000$. The top and bottom column show average interval volume and cumulative miscoverage rate over $1000$ repetitions with $\mathrm{df}=\infty$.
S7.1.3 Multivariate biased predictions
Here, we illustrate the multivariate AsympCS procedure described in Appendix˜S4 in the context of PPI. To do this, we study a simple multivariate version of the mean estimation task with biased prediction described in Section˜6.1. In particular, for $d=5$, we sample $d$-dimensional observations $Y_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mathbf{0},\Sigma)$, where $\Sigma\in\mathbb{R}^{d\times d}$ is a Toeplitz covariance matrix with entries $\Sigma_{ij}=0.5^{|i-j|}$, and define $Y_{i}=f(X_{i})+\epsilon_{i}$, where $\epsilon_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mathbf{0},I_{d})$, so that $\theta^{\star}=\mathbb{E}[Y]=\mathbf{0}$. Then, we proceed as in Section˜6.1 and define biased predictions $f(X_{i})=X_{i}+\upsilon$, where $\upsilon\in\mathbb{R}$ controls the bias level of the predictor. In this setup, we compare classical inference, non-assisted PPI, and Bayes-assisted PPI under a Gaussian prior with mean zero and isotropic covariance matrix using the spherical multivariate AsympCS procedures described in Appendix˜S4. The AsympCS hyperparameters are set using natural extensions of the rules in Section˜S6.4 to the multivariate case, with $t^{\star}=1000$ for all methods. Figure˜2 shows the average spherical interval volumes of each method as a function of $\upsilon$, which we vary between 6-6 and $6$, at $n\in\{100,250,500\}$.
Figure S9: Multivariate biased predictions study. The left, middle and right panels show average spherical region radius over $1000$ repetitions for $n=100,250,500$.
The results are consistent with those in Figure˜2. In particular, non-assisted PPI consistently outperforms classical inference, with both methods yielding constant interval radii across bias levels. On the other hand, Bayes-assisted PPI achieves smaller radii than the other baselines for small values of $\upsilon$, but its radius grows quickly with $|\upsilon|$ as the prior becomes increasingly misspecified. For this example, we do not report coverage results, as we find that all methods achieve near perfect coverage across the values of $\upsilon$ considered, with cumulative miscoverage rates close to zero, likely due to the conservative spherical construction mentioned in Appendix˜S4.
S7.2 Real data
S7.2.1 Mean estimation
@@
For each of the mean estimation experiments, we set $t^{\star}$ (see Section˜S6.4) equal to the largest $n$ considered in the experiment. In particular, for the flights, forest, and galaxies datasets, we set $t^{\star}=10000$, $500$, and $1000$, respectively.
Figure˜S10 adds the results of the standard PPI procedures to the ones shown in Figure˜3.
@@
Figure S10: Mean estimation. The top and bottom rows show the average interval volume and cumulative miscoverage rate over $1000$ repetitions for the flights, forest, and galaxies datasets.
@@
For these experiments, the improvement of PPI++ over standard PPI is small, and the results remain consistent with those in Figure˜3. That is, PPI methods consistently improve over classical inference, with Bayes-assisted methods providing an additional efficiency boost for moderate labelled sample sizes.
S7.2.2 Other estimation tasks
The estimation tasks considered here involve linear regression (census dataset), logistic regression (healthcare dataset), and median estimation (genes dataset). As above, for each estimation task, we set $t^{\star}=2000$ (see Section˜S6.4), as that is the largest $n$ considered in all experiments.
For these, AsympCS procedures relying on classical inference (obtained from Theorem˜1) and PPI, both non-assisted (Proposition˜3) and Bayes-assisted (Equation˜26), require constructing a grid over $\theta$ through Equation˜4. To initialise the grid, we use the first $n_{0}$ labelled data points to compute a preliminary estimate of $\theta^{\star}$, which we then use to centre the grid. The same $n_{0}$ is also used as the starting point to evaluate the AsympCS procedures and compute their cumulative miscoverage rate reported in the figures below. We set $n_{0}=100$ for the census and healthcare datasets, and $n_{0}=40$ for the genes dataset.
Furthermore, some priors, including the Student-$t$ prior, require numerical integration to compute the marginal density $\eta_{t}$ used in Equation˜26. As a result, when Bayes-assisted PPI under such priors is used, the computational cost grows significantly when Equation˜26 is evaluated across many $n$ and $\theta$ values simultaneously. Because of this, we only report results for the Gaussian and Laplace priors, which admit closed-form expressions for $\eta_{t}$.
Figure˜S11 compares classical and PPI AsympCS procedures on the three estimations tasks above in terms of average interval volume and cumulative miscoverage rate as $n$ increases.
Figure S11: Other estimation tasks. The top and bottom rows show the average interval volume and cumulative miscoverage rate over $100$ repetitions for the census, healthcare, and genes datasets.
As discussed in Section˜6.2, PPI methods outperform classical inference for the linear and logistic regression tasks, with Bayes-assisted methods further improving efficiency when $n$ is moderate. For the median estimation task, on the other hand, non-assisted PPI still improves over classical inference, while Bayes-assisted PPI yield larger regions than the other methods due to the higher bias of the predictions in this dataset. In all cases, coverage remains satisfactory.
As discussed in Section˜S6.2, the census, healthcare, and genes datasets are preprocessed by applying a log-transformation to relevant positive skewed variables, as it is commonly done in the literature. For instance, this is the case for the income variable $Y_{i}$ in the census dataset. In practice, such a transformation improves the accuracy of the KMT coupling approximation for a given $n$, essentially lowering the effective labelled sample size necessary to achieve satisfactory coverage. To see this, we repeat the linear regression experiment on the census dataset without applying any preprocessing to $Y_{i}$. As shown in Figure˜S12, the results are strikingly different: all methods yield significantly larger cumulative miscoverage rates compared to the preprocessed case in Figure˜S11, with non-assisted PPI violating the nominal guarantee around $n\approx 500$, in turn invalidating the efficiency comparison.
Figure S12: Linear regression on the census dataset without preprocessing. The top and bottom rows show the average interval volume and cumulative miscoverage rate over $100$ repetitions.
Without preprocessing, a substantially larger starting labelled sample size $n_{0}$ is needed before the KMT coupling approximation is accurate enough for satisfactory uniform-time coverage by the AsympCS procedures. This example highlights the importance of knowledge of the data distribution when using AsympCS procedures in practice. A possible way to obtain such knowledge in practice is to estimate the third moment of the data distribution, if it exists, from a held-out validation set and use a Berry-Esseen-type bound [28, Theorem 2.1.4] to choose a starting labelled sample size $n_{0}$ at which the KMT coupling provides a good approximation.
Appendix S8 Alternative non-assisted AsympCS
S8.1 Parameter-free AsympCS via improper prior
As discussed in Section˜3.1, the non-assisted asymptotic confidence sequence $\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t};\rho)$ in Equation˜6 approximates the exact CS in Equation˜S30 and becomes arbitrarily accurate as in the limit. This suggests constructing alternative non-assisted AsympCS by approximating other exact CSs for which adaptations of Theorems˜1 and 2 apply.
One example is the parameter-free non-assisted CS of Wang and Ramdas [19, Corollary 5.9]. Define the continuous, strictly decreasing bijection $g\colon[1,\infty)\to(0,1]$ by
$g(x):=2\left[1-\Phi\left(\sqrt{\log(x^{2})}\right)\right]+2\sqrt{\log(x^{2})}\phi\left(\sqrt{\log(x^{2})}\right),$
with $\Phi$ and $\phi$ the standard normal CDF and PDF, and let $z_{\alpha}:=g^{-1}(\alpha)$, which is well-defined. The corresponding AsympCS is given by
$\mathcal{C}^{\mathrm{\tt NA}^{\prime}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t}):=\left[\overline{Y}_{t}\pm\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(tz_{\alpha}^{2}\right)}\right].$ (S90)
Notably, for i.i.d. Gaussian observations with known variance, the exact (nonasymptotic) counterpart of (S90) is obtained by applying the method of mixtures for extended nonnegative martingales [19, Def. 3.1] together with extended Ville’s inequality [19, Theorem 4.1], using a non-informative improper prior as the mixing density.
S8.2 Experiments
We use $\mathcal{C}^{\mathrm{\tt NA}^{\prime}}_{\alpha,t}$ from (S90) as a parameter-free drop-in replacement for $\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}$ in the classical and prediction-powered AsympCS procedures from the main text and repeat some of the experiments from Section˜6. In figures, runs that use the alternative AsympCS are annotated (I) for “improper”.
@@
Figure S13: Noisy predictions study with alternative AsympCS . The left, middle and right panels show average interval volume and cumulative miscoverage rate over $1000$ repetitions for noise levels $\sigma_{Y}=0.1,0.8,3.0$. Results for non-assisted PPI++ based on Equation˜6 are shown for reference.
@@
Figure S14: Biased predictions study with alternative AsympCS . The left, middle and right panels show average interval volume and cumulative miscoverage rate over $100$ repetitions for $\mathrm{df}=5,10,\infty$. Results for non-assisted PPI based on Equation˜6 are shown for reference.
@@
Figure S15: Real data study with alternative AsympCS . The top and bottom rows show the average interval volume and cumulative miscoverage rate over $1000$ repetitions for the flights, forest, and galaxies datasets. Results for classical inference and non-assisted PPI++ based on Equation˜6 are shown for reference.
Compared with the standard non-assisted AsympCS used in the main text, which depends on the hyperparameter $\rho$, the parameter-free alternative typically performs slightly worse under our default choice of $\rho$ (see Section˜S6.4). Nonetheless, $\mathcal{C}^{\mathrm{\tt NA}^{\prime}}_{\alpha,t}$ can represent an attractive choice when selecting $\rho$ is problematic, precisely because it avoids any tuning.